OpenStudy (anonymous):

How do I simplify this? (2√5+3√7)^2

OpenStudy (cicilybailey14):

(2√5+3√7)2 =((2)(2.236068)+3√7)2 =(4.472136+3√7)2 =(4.472136+(3)(2.645751))2 =(4.472136+7.937254)2 =12.409392 Answer =153.992957

OpenStudy (anonymous):

:O Thanks I was wanting for you explain but its ok thank you so much @Cicilybailey14 :)

OpenStudy (cicilybailey14):

Oh im sorry! and your wlecome :)